a. What is the net electric flux through the surface is it encloses a net charge of 3.60μC?


b. The electric flux through a closed surface is found to be 7.80Nm2


/C. What quantity


of charge is closed by the surface?



1
Expert's answer
2022-04-22T13:53:20-0400

(a)

"\phi=\frac{q}{\epsilon_0}=\frac{3.60*10^{-6}}{8.85*10^{-12}}=4.07*10^5\:\rm N\cdot m^2/C"

(b)

"q=\phi\epsilon_0=7.80*8.85*10^{-12}=6.90*10^{-11}\:\rm C"